# A study of gamma-ray emission from OJ 287 using Fermi-LAT from 2015-2023

### v1

@@ 3.9 Segment 8 (JD: 2459810-2459999)
The photon flux light curve and SED for segment 8 can be found in Fig. 25 and  26, respectively. We note that the optical data was found to be in a low flux state in  Gupta et al. (2023). The average photon and energy flux are given by $(4.61\pm 0.18)\times 10^{-8}\rm{cm^{-2}s^{-1}}$ and $(2.15\pm 0.14)\times 10^{-5}\rm{MeV~{}cm^{-2}s^{-1}}$ respectively. We find that only six data points in the light curve have non-zero values, with the rest showing upper limits or flux values consistent with zero. Similarly the SED only shows values until 2 GeV. The posteriors for the best-fit spectral parameters can be found in Fig. 27.
Figure 25: Segment 8 (MJD : 59809.5 - 59998.5 ) - Light Curve
Figure 26: Segment 8 - SED
Figure 27: Segment 8 - SED Best fit parameters: $\log_{10}N_{0}=-10.07\pm 0.07$ ; $\alpha=1.96\pm 0.08$ ; $E_{0}=705.82$ MeV
@@ 3.10 Segments 1 - 8 (JD: 2457284 - 2459999)
We now do a holistic analysis of the data for all the eight segments. The photon flux light curve and SED for segment 8 can be found in Fig. 28 and  29, respectively. The average photon and energy flux are given by $(5.16\pm 0.05)\times 10^{-8}\rm{cm^{-2}s^{-1}}$ and $(2.71\pm 0.04)\times 10^{-5}\rm{MeV~{}cm^{-2}s^{-1}}$ respectively. From the SED, we find non-zero flux up to approximately 50 GeV and upper limits beyond that. Similar to segment 3A, we see an approximately constant flux, except for the first data point. We see a falling light curve from the start of segment 1 unil MJD of around 58000. We then see a quasi-sinusoidal trend in the light curves. The posteriors for the best-fit spectral parameters can be found in Fig. 30. The isotropic average gamma-ray luminosity detected during the entire segment is $2.18\times 10^{47}$ ergs/sec.
Figure 28: Segments 1-8 (MJD : 57283.5 - 59998.5 ) - Light Curve
Figure 29: Segments 1-8 - SED
Figure 30: Segments 1-8 - SED Best fit parameters: $\log_{10}N_{0}=-9.87\pm 0.01$ ; $\alpha=2.12\pm 0.02$ ; $E_{0}=715.56$ MeV
Segment TS Flux ×$\times 10^{-8}$ Energy Flux ×$\times 10^{-5}$ $\log(N0)$ $\alpha$ Luminosity ×$\times 10^{47}$ Pivot Energy
(cm-2 s-1) (MeV cm-2 s-1) (erg s-1) $E_{0}$ (MeV)
1 1586.61 $10.79\pm 0.70$ $4.65\pm 0.27$ $-9.51\pm 0.03$ $2.14\pm 0.04$ 3.78 636.6
2 891.89 $4.07\pm 0.05$ $3.52\pm 0.09$ $-9.94\pm 0.05$ $1.92\pm 0.05$ 2.76 720.59
3A 155.29 $3.44\pm 0.95$ $1.53\pm 0.26$ $-9.89\pm 0.16$ $2.19\pm 0.19$ 2.76 846.77
3B 154.6 $3.88\pm 0.67$ $1.86\pm 0.33$ $-10.06\pm 0.12$ $1.97\pm 0.12$ 2.31 803.72
4 263.04 $4.80\pm 0.78$ $2.02\pm 0.34$ $-9.91\pm 0.06$ $2.21\pm 0.10$ 0.98 545.92
5 379.65 $4.52\pm 0.72$ $2.60\pm 0.45$ $-9.96\pm 0.05$ $2.10\pm 0.07$ 3.1 921.45
6 308.48 $3.72\pm 0.67$ $2.06\pm 0.33$ $-9.97\pm 0.06$ $2.03\pm 0.08$ 2.18 745.69
7 320.09 $2.06\pm 0.25$ $1.95\pm 0.27$ $-10.20\pm 0.13$ $1.84\pm 0.12$ 3.96 1081.02
8 173.44 $4.61\pm 0.18$ $2.15\pm 0.14$ $-10.07\pm 0.07$ $1.96\pm 0.08$ 1.79 705.82
all 4995.36 $5.16\pm 0.05$ $2.71\pm 0.04$ $-9.87\pm 0.01$ $2.12\pm 0.02$ 2.18 715.56
@@
Table 1: Summary of test statistics (TS), observed photon and energy flux, pivot energy, $E_{0}$ (cf. Eq 1) along with the best-fit SED parameters for all the eight observing segments. We note that the detection significance is given by the square root of TS (Mattox et al., 1996).
3.11 Comparison with Prince et al
@@
We now compare our results with those in  Prince et al. (2021a). A summary of their observations can be found in Table 4. We find that the spectral index $\alpha$ is in the same ball-park range between 1.85 and 2.24 found in  Prince et al. (2021a). The photon flux in segment 1 is about a factor of two larger than the fluxes detected in Prince et al. (2021a). For all other segments it is comparable. The isotropic luminosity that we detect, which is between $10^{47}-10^{48}$ ergs/sec, is comparable to that detected in  Prince et al. (2021a) and is less than the Eddington luminosity of $10^{50}$ erg/sec. Therefore, our results are mostly consistent with those in  Prince et al. (2021a).
4 Conclusions
@@
In a recent work, Gupta et al. (2023), presented the results of an extensive observing campaign of OJ 287 from 2015-2023 using optical telescopes in three different continents in conjunction with public archival data from the Steward Observatory. These observations were divided into eight segments, where each segment roughly corresponds to an observing season. The third segment was further sub-divided into two sub-segments. We searched for coincident gamma-ray emission from 0.1-300 GeV with the Fermi-LAT telescope during each of the observing segments. For each segment, we show the gamma-ray light curve and SED. We fit the SED to a power law and show a corner plot depicting the best-fit parameters. We also calculate the luminosity for each segment. the plots for the gamma-ray light curves, SED, and marglinalized contours for the SED best-fit parameters for all the eight segments can be found in Figs. 1-30. A tabular summary of our results can be found in Table 1. Our conclusions are as follows:
@@
  • The observed energy flux is between $(2-4.6)\times 10^{-5}MeVcm^{-2}s^{-1}$.
@@
  • The maximum flux is seen near the start of first segment. The flux decreases unil MJD of 58000.
  • We do not find any correlated gamma-ray flare during Segment 7 in coincidence with an optical flare in $I$ and $R$ band recently reported in  Valtonen et al. (2024).
  • We do not detect any emission beyond 20-30 GeV.
@@ Acknowledgments
We are grateful to Raniere Menezes for help with the EasyFermi package.

### v2

@@ 3.9 Segment 8 (JD: 2459810-2459999)
The photon flux light curve and SED for segment 8 can be found in Fig. 25 and  26, respectively. We note that the optical data was found to be in a low flux state in  Gupta et al. (2023). The average photon and energy flux are given by $(4.61\pm 0.18)\times 10^{-8}\rm{cm^{-2}s^{-1}}$ and $(2.15\pm 0.14)\times 10^{-5}\rm{MeV~{}cm^{-2}s^{-1}}$ respectively. We find that only six data points in the light curve have non-zero values, with the rest showing upper limits. Similarly, the SED only shows values until 2 GeV. The posteriors for the best-fit spectral parameters can be found in Fig. 27.
Figure 25: Segment 8 (MJD : 59809.5 - 59998.5 ) - Light Curve ; binwidth = 9.45 days.
Figure 26: Segment 8 - SED. The black solid line shows the power-law fit while the gray line shows the log-parabola fit from Fermipy.
Figure 27: Segment 8 - SED Best fit parameters: $\log_{10}N_{0}=-10.07\pm 0.07$ ; $\alpha=1.96\pm 0.08$ ; $E_{0}=705.82$ MeV.
@@ 3.10 Segments 1 - 8 (JD: 2457284 - 2459999)
We now do a holistic analysis of the entire data for all eight segments. The photon flux light curve and SED for segment 8 can be found in Fig. 28 and  29, respectively. The average photon and energy flux are given by $(5.16\pm 0.05)\times 10^{-8}\rm{cm^{-2}s^{-1}}$ and $(2.71\pm 0.04)\times 10^{-5}\rm{MeV~{}cm^{-2}s^{-1}}$ respectively. From the SED, we find non-zero flux up to approximately 20 GeV and upper limits beyond that. Similar to segment 3A, we see an approximately constant flux, except for the first data point. We see a falling light curve from the start of segment 1 until MJD of around 58000. We also confirmed that we do not find any signal beyond 20 GeV even aftercoarser binning. The posteriors for the best-fit spectral parameters can be found in Fig. 30. The isotropic average gamma-ray luminosity detected during the entire segment is $2.18\times 10^{47}$ ergs/sec.
We also did a search for periodicity using the data for all eight segments using the generalized Lomb-Scargle periodogram (VanderPlas, 2018). For this purpose, we use the same implementation as our most recent work on the analysis of Super-Kamiokande solar neutrino data (Pasumarti and Desai, 2024). We searched using a frequency grid with a resolution of 0.028 $yr^{-1}$ and ranging from 0 to 2.69 $yr^{-1}$. The largest LS power is at 2.15 $yr^{-1}$. However, the $p$-value using all the metrics specified in  VanderPlas (2018) is greater than 0.1. Therefore, we conclude that there is no discernible periodicity in the OJ 287 gamma-ray dataset during the above period.
Figure 28: Segments 1-8 (MJD : 57283.5 - 59998.5) - Light Curve ; binwidth = 135.75 days.
Figure 29: Segments 1-8 - SED. The black solid line shows the power-law fit while the gray line shows the log-parabola fit from Fermipy.
Figure 30: Segments 1-8 - SED Best fit parameters: $\log_{10}N_{0}=-9.87\pm 0.01$ ; $\alpha=2.12\pm 0.02$ ; $E_{0}=715.56$ MeV.
Segment TS Flux ×$\times 10^{-8}$ Energy Flux ×$\times 10^{-5}$ Luminosity
(cm-2 s-1) (MeV cm-2 s-1) ×$\times 10^{47}$ (erg s-1)
1 1586.61 $10.79\pm 0.70$ $4.65\pm 0.27$ 3.78
2 891.89 $4.07\pm 0.05$ $3.52\pm 0.09$ 2.76
3A 155.29 $3.44\pm 0.95$ $1.53\pm 0.26$ 2.76
3B 154.6 $3.88\pm 0.67$ $1.86\pm 0.33$ 2.31
4 263.04 $4.80\pm 0.78$ $2.02\pm 0.34$ 0.98
5 379.65 $4.52\pm 0.72$ $2.60\pm 0.45$ 3.1
6 308.48 $3.72\pm 0.67$ $2.06\pm 0.33$ 2.18
7 320.09 $2.06\pm 0.25$ $1.95\pm 0.27$ 3.96
8 173.44 $4.61\pm 0.18$ $2.15\pm 0.14$ 1.79
all 4995.36 $5.16\pm 0.05$ $2.71\pm 0.04$ 2.18
@@
Table 1: Summary of test statistics (TS), observed photon, energy flux, and the isotropic gamma-ray luminosity. We note that the detection significance is given by the square root of TS (Mattox et al., 1996).
Simple PL fit Log Parabola fit
Segment Pivot $E_{0}$ (MeV) $\log_{10}$0 $\alpha$ $\log_{10}$0 $\alpha$ $\beta$ $E_{b}$ (MeV)
1 $-9.51\pm 0.03$ $2.14\pm 0.04$ 636.6 $-10.65\pm 0.04$ $2.15\pm 0.04$ $0.10\pm 0.02$ 710.99
2 $-9.94\pm 0.05$ $1.92\pm 0.05$ 720.59 $-10.97\pm 0.01$ $1.80\pm 0.01$ $0.07\pm 0.00$ 710.99
3A $-9.89\pm 0.16$ $2.19\pm 0.19$ 846.77 $-11.10\pm 0.13$ $2.11\pm 0.16$ $0.14\pm 0.08$ 710.99
3B $-10.06\pm 0.12$ $1.97\pm 0.12$ 803.72 $-11.16\pm 0.10$ $2.15\pm 0.09$ $0.03\pm 0.04$ 710.99
4 $-9.91\pm 0.06$ $2.21\pm 0.10$ 545.92 $-11.15\pm 0.12$ $2.27\pm 0.09$ $0.00\pm 0.04$ 710.99
5 $-9.96\pm 0.05$ $2.10\pm 0.07$ 921.45 $-11.11\pm 0.09$ $2.13\pm 0.08$ $0.00\pm 0.03$ 710.99
6 $-9.97\pm 0.06$ $2.03\pm 0.08$ 745.69 $-11.11\pm 0.11$ $2.04\pm 0.10$ $0.05\pm 0.05$ 710.99
7 $-10.20\pm 0.13$ $1.84\pm 0.12$ 1081.02 $-11.19\pm 0.09$ $1.69\pm 0.07$ $0.12\pm 0.03$ 710.99
8 $-10.07\pm 0.07$ $1.96\pm 0.08$ 705.82 $-11.26\pm 0.03$ $2.33\pm 0.02$ $-0.07\pm 0.01$ 710.99
all $-9.87\pm 0.01$ $2.12\pm 0.02$ 715.56 $-11.02\pm 0.01$ $2.11\pm 0.01$ $0.03\pm 0.00$ 710.99
Table 2: Results of spectral fits using the power-law (cf. Eq. 1) as well as log-parabola model (cf. Eq. 2) for all the eight segments
4 Comparison with other monitoring programs for OJ 287
4.1 Comparison with Prince et al.
@@ 4 Comparison with other monitoring programs for OJ 287
We now compare our results with the gamma-ray observations for OJ 287 reported in  Prince et al. (2021a). A summary of their observations can be found in Table 4. We find that the spectral index $\alpha$ is in the same ball-park range between 1.85 and 2.24 found in  Prince et al. (2021a). The photon flux in segment 1 is about a factor of two larger than the fluxes detected in Prince et al. (2021a). For all other segments, it is comparable. The isotropic luminosity that we detect, which is between $10^{47}-10^{48}$ ergs/sec, is comparable to that detected in  Prince et al. (2021a) and is less than the Eddington luminosity of $10^{50}$ erg/sec. Therefore, our results are mostly consistent with those in  Prince et al. (2021a).
Figure 31: Comparison of the optical $R$-band observations reported in  Valtonen et al. (2024) alongside the flux observed by Fermi-LAT during Segment 7. The gamma-ray and optical light curves have been normalized by their respective modes.
4.2 Comparison with $R$-band optical flare in 2021-2022
Most recently,  Valtonen et al. (2024) reported optical flares in the $R$-band between October 21st and November 2021 as part of the Krakow quasar monitoring campaign consisting of several telescopes located all over the globe. These flares are caused by the impact of the secondary black hole onto the accretion disk of the primary black hole. This flare occurred during segment 7. To compare the optical data with the Fermi-LAT data, we plot both in the same figure after normalizing each data by the mode. This plot can be found in Fig. 31. As we can see, although the gamma-ray light curve is sampled coarsely compared to the optical light curve, there is no direct one-to-one correlation between the rise and fall time of the gamma-ray and optical light curves during this segment. The fall in the optical light curve coincides with the rising gamma-ray flux at the beginning of the segment, where the peaks in the two light curves are separated by around 200 days.
4.3 Comparison with SWIFT XRT light curves from 2015-2021
OJ 287 has also been monitored by the SWIFT satellite in optical/UV bands using six filters as well as X-rays from 0.3-10 keV as part of the MOMO (Multiwavelength Observations and Modelling of OJ 287) project (Komossa et al., 2023a) from September 2015 till September 2021. Additional SWIFT observations starting from 2015 have also been reported (Komossa et al., 2020, 2021). These observations have been collated in  Komossa et al. (2022). Theoretical implications of the observations from the MOMO project can be found in  Komossa et al. (2023b, c). We overlay these SWIFT X-ray observations along with the Fermi-LAT observations in Fig. 32. Once again we do not see any correlation between the peaks in the SWIFT XRT light curves and gamma-ray light curves.
Figure 32: Comparison of the SWIFT XRT observations from 0.3-10 keV obtained from Komossa et al. (2022), alongside the flux observed by Fermi-LAT. The gamma-ray and X-ray observations are normalized by their respective modes.
5 Discussion and Conclusions
@@
In a recent work, Gupta et al. (2023) presented the results of an extensive observing campaign of OJ 287 from 2015-2023 using optical telescopes in three different continents in conjunction with public archival data from the Steward Observatory. These observations were divided into eight segments, each roughly corresponding to an observing season. The third segment was further subdivided into two sub-segments. We searched for coincident gamma-ray emission from 0.1-300 GeV with the Fermi-LAT telescope during each observing segment. For each segment, we show the gamma-ray light curve and SED. We fit the SED to a power law and show a corner plot depicting the best-fit parameters. We also calculate the luminosity for each segment. the plots for the gamma-ray light curves, SED, and marginalized contours for the SED best-fit parameters for all the eight segments can be found in Figs. 1-30. Tables  1 and  2 provides a tabular summary of our results. Our conclusions are as follows:
@@
  • The observed energy flux is between $(2-4.6)\times 10^{-5}~{}\rm{MeVcm^{-2}s^{-1}}$.
@@
  • The maximum flux is seen near the start of the first segment. The flux decreases until MJD of 58000.
  • We do not find any correlated gamma-ray flare during Segment 7 in coincidence with an optical flare in $I$ and $R$ band recently reported in  Valtonen et al. (2024). A comparison of the gamma-ray and optical light curves can be found in Fig. 31.
  • We also do not find any correlation between the peaks in the gamma-ray flux and X-ray flux observed by the SWIFT satellite between 0.3-10 keV (cf. Fig 32).
  • We do not detect any emission beyond 20-30 GeV, similar to previously reported gamma-ray observation from OJ287 (Prince et al., 2021a; Kushwaha et al., 2013). The gamma-ray spectrum in OJ287 has been previously explained due to a combination of Synchrotron self-Compton and external Compton processes from the scattering of soft photons that are outside the jet (Kushwaha et al., 2013). The cut-off in the spectrum can be explained with a power law distribution of electrons with a maximum energy of approximately $10^{4}$ times the electron rest mass Kushwaha et al. (2013).
@@ Acknowledgments
We are grateful to Raniere Menezes for his help with the EasyFermi package. We are grateful to Stefanie Komossa for providing us the SWIFT X-ray data from  Komossa et al. (2022) and Mauri Valotnen for $R$-band data reported in  Valtonen et al. (2024). We also acknowledge the anonymous referee for constructive feedback on our manuscript.